# A resource-efficient variational quantum algorithm for mRNA codon optimization

### v1

@@ 1 Introduction
In bioinformatics, computational problems often present significant challenges due to the massive amount of data and the substantial computational resources. Searching for effective solutions to these problems is important for the development of bioinformatics. Quantum computing emerges as a promising candidate due to its distinctive properties, including superposition, entanglement and interference between computational paths. These properties enable quantum computers to process vast amounts of data in parallel, potentially overcoming the limitations of classical computing for bioinformatics tasks. Therefore, exploring quantum algorithms for bioinformatics holds significant promise for accelerating the advancements in this field.
@@ 2 Related research
In the realms of DNA or RNA vaccine development, bioinformatics leverages the analysis of extensive pathogen sequences to identify target antigenic proteins for mRNA sequence design. This expedites the vaccine design process, facilitating the creation of safe and effective vaccines tailored to combat specific viruses. The COVID-19 pandemic, which spread rapidly from 2019 to 2022, had a profound global impact, affecting a vast number of people. This surge strained healthcare systems, leading to a scarcity of health resources and thereby preventing many individuals from receiving timely treatment. This had wide-ranging consequences on global health and everyday life. Additionally, the pandemic triggered a surge in unemployment, trade disruptions, and a downturn in the global economy. [4] This underscored the urgent need to conduct research into vaccine development, particularly focusing on highly infectious viruses, to safeguard human health.
@@
  • mRNA vaccines - mRNA vaccines have emerged as a revolutionary approach to immunization, utilizing a fragment of messenger RNA (mRNA) to trigger the production of viral proteins (antigens) within the human body. These antigens, produced by human cells in response to the injected mRNA, stimulate the immune system to generate protective antibodies. This cutting-edge technique not only offers safety advantages by not using weakened or dead viruses but also exhibits rapid development, high protein expression efficiency, and shortened production timelines. [5][7][8]
  • Structure of the mRNA sequence - As shown in Figure 2, the structure of mRNA sequences is composed of the following components: 5’ cap, 5’ Untranslated Region(5’ UTR), Open Reading Frame(ORF), 3’ Untranslated Region (UTR), and Ploy-A tail. [9] While the 5’ untranslated region (5’UTR) and the 3’ untranslated region (3’UTR) are essential in regulating protein expression, they do not directly determine the specific amino acids incorporated into the protein. It’s worth noting that this paper does not discuss the optimization of mRNA codons through the 5’UTR and 3’UTR regions. The ORF is a part of the DNA, RNA, or mRNA sequence, which begins with a start codon (usually AUG, CUG, UUG in mRNA) and ends with a stop codon(UAA, UAG, and UGA in mRNA). The ORF in mRNA serves as the blueprint for protein synthesis, dictating the specific sequence of amino acids to be incorporated with the resulting antigen protein. In addition, the ORF is crucial for the development of mRNA vaccines, as it also plays a pivotal role in protein expression and stability.[10][11]
@@
    Traditional methods for solving this problem on classical computers can be computationally intensive, especially as the length of the target sequence increases. The complexity arises from the vast search space of possible codon combinations and the need to balance factors like codon usage bias and secondary mRNA structure. Deep learning, particularly neural networks, has shown promise in addressing complex bioinformatics problems, including mRNA codon optimization. [13][14][15] Neural networks can automatically learn meaningful representations of biological data. In the context of mRNA codon optimization, this means that the model can capture the essential features of codon sequences that lead to optimal protein expression. Pre-trained models on related tasks or datasets can be fine-tuned for specific mRNA codon optimization problems. This can leverage knowledge gained from one domain to enhance performance in another. While deep learning holds promise, it’s essential to note that the success of these models depends on the availability of high-quality training data and careful design of the neural network architecture. Additionally, interoperability and understanding the biological implications of the model’s predictions are crucial in the context of bioinformatics applications.
@@
      Quantum annealers, like those developed by D-Wave Systems, hold significant promise for solving combinatorial optimization problems. [20][21][22] Quantum annealing is an area of research with practical applications, including in the field of mRNA codon optimization. [3] The general process of applying quantum annealer to combinatorial optimization problems, such as mRNA codon optimization involves several steps. One is problem definition, specifying the optimization goal and relevant constraints. The second is to transform the problem into an Ising model representation. Then, the quantum annealing devices are used to find the lowest energy state. The final step is to decode the result to obtain the mRNA codon sequence, reflecting an optimization solution for protein expression.
@@
      VQE is a quantum algorithm designed to find the minimum eigenvalue for a given Hamiltonian, making it well-suited for solving certain types of combinatorial problems.[23] Firstly, a combinatorial problem is defined using a Hamiltonian. Then, a circuit ansatz with variational parameters is optimized to the approximate ground state of the Hamiltonian corresponding to the minimum expectation value. These quantum circuits are executed on a quantum computer to obtain the expectation value of the Hamiltonian. The process is iterated until convergence or a satisfactory solution quality is reached.
@@
      QAOA is a variational quantum algorithm, belonging to a broader class of VQE. It is tailored to find approximate optimization problems’ solutions. [24] The algorithm involves mapping a specific problem to the Ising model. In the aspect of quantum circuit construction, quantum state preparation encodes potential solutions, and a quantum circuit with alternating layers of unitary operations is built. Parameters are adjusted using classical optimization techniques to minimize the expectation value of the Hamiltonian. Iterations continue until convergence is reached.
@@
We have developed an innovative quantum method that significantly reduced quantum computing resources with a dense way of encoding compared to the existing quantum approach. [3] However, our approach adds complexities to the Hamiltonian due to the multiple local terms. Quadratic unconstrained binary optimization (QUBO) and QAOA approach mandates a Hamiltonian composed of only two local terms and it needs to add extra qubits for the reduction of additional terms. But this reduction also results in an increase in the required number of qubits. To address these complexities, we have employed the VQE algorithm, as it offers a unique advantage: it does not demand the addition of qubits. VQE enables us to seek the ground energy of the objective function efficiently, following which we employ classical computing techniques to derive the optimal mRNA sequence based on the ground state.
@@
Qiskit platform – Our implementation of codon optimization is conducted using the Qiskit platform which is the gate circuit quantum simulator as a central tool for codon optimization. Nevertheless, current quantum computing resources are constrained by the number of qubits. For example, optimizing the full-length protein sequences would demand thousands of qubits, surpassing the capabilities of existing quantum hardware. We’ve adopted a pragmatic approach to overcome this limitation by dividing the protein sequence into smaller, more manageable fragments with shorter lengths. This strategic division enables us to operate with fewer qubits for each fragment, ensuring a more feasible and efficient optimization process. By presenting this innovative approach to mRNA codon optimization, we aim to contribute to the evolving field of quantum computing in bioinformatics, providing new perspectives and solutions to complex problems in the realm of genetic coding and sequence optimization. The Qiskit platform allows to execute quantum circuits in two ways. The first one connects to a native QPU backend like superconducting qubits, taking the results from the latest quantum processors into account. The second one is using quantum circuit simulator where the qubits are assumed to be perfect qubits. That intrinsically implies no decoherence of the qubits and that the quantum gates have no error. Error models can be applied on the simulator for realistic simulation, however, often becomes a barrier for quantum logic and algorithm development for future quantum applications. For this project, we employ noiseless quantum circuit simulation for the proof-of-concept development and demonstration of our algorithm, based on the Perfect Intermediate Scale Quantum (PISQ) [25] approach.
@@ 3 Method
Quantum computers, in particular, hold promise due to their distinctive quantum mechanical properties such as superposition, entanglement and interference. These properties empower quantum computers to efficiently handle extensive quantum data sets, finding the optimal (lowest or highest) value in a vast space solution. Notably, Fox et al. [3] introduced a codon optimization algorithm implemented on quantum annealers and gate-based circuit quantum computers. This groundbreaking work offers a practical approach to mRNA sequence design, unlocking new possibilities in the realm of bioinformatics and genetic research. The research by these authors represents a crucial milestone in harnessing the capabilities of quantum computing for this innovative application.
@@ 3.1 Synonymous codons encoding
Two examples for encoding amino acids A and L are given in Table 2 using either the one-hot or the dense way. This encoding allows straightforward identification of the selected codon based on the binary pattern. However, it introduces redundant quantum states (explained in Appendix), which in turn increases the complexity of the quantum system.
@@ 3.3 Quantum algorithm execution
The qasm-simulator on the Qiskit platform was used to execute the VQE algorithm for the mRNA codon optimization for the protein fragment with a length of 8 while utilizing between 8 and 21 qubits. Given the challenges in accessing real quantum computers with a sufficient number of qubits, simulators in Qiskit allow scientists to explore optimization problems in quantum computing. In addition, Python3 served to generate Hamiltonian and as an accessible programming language on Qiskit. In the final step, the outcome of Qiskit was translated to mRNA sequence through Python3.
@@ 4 Results
We selected the protein P0DTC2 for optimizing mRNA codon sequences. The P0DTC2 protein from the spike protein of 2019-nCoV, sourced from UniProt and consisting of 1273 amino acids, would require 2234 and 4417 qubits for dense and one-hot encoding respectively according to equations 1 and 2. However, the available number of qubits on current quantum computers is limited, with no quantum computer possessing more than 1000 qubits. It is unrealistic to achieve mRNA codon optimization for the full-length protein P0DTC2 on a quantum computer. As a result, we divided the protein into 159 fragments, each with a length of 8 amino acids (excluding the first amino acid M, which has only one codon and does not need optimization). Consequently, the number of required qubits for each fragment falls within the range of 8 to 21, as determined by equation 2. Each fragment is used to construct a Hamiltonian via equation 7, and then VQE is employed to identify the optimal codon sequence for each fragment. We selected the protein P0DTC2 for optimizing mRNA codon sequences. The P0DTC2 protein from the spike protein of 2019-nCoV, sourced from UniProt and consisting of 1273 amino acids, would require 2234 and 4417 qubits for dense and one-hot encoding respectively according to equations 1 and 2. However, the available number of qubits on current quantum computers is limited, with no quantum computer possessing more than 1000 qubits. It is unrealistic to achieve mRNA codon optimization for the full-length protein P0DTC2 on a quantum computer. As a result, we divided the protein into 159 fragments, each with a length of 8 amino acids (excluding the first amino acid M, which has only one codon and does not need optimization). Consequently, the number of required qubits for each fragment falls within the range of 8 to 21, as determined by equation 2. Each fragment is used to construct a Hamiltonian via equation 7, and then VQE is employed to identify the optimal codon sequence for each fragment. We use paper [3] as our reference paper.
(a) The required number of qubits distribution between one hot and dense encoding on proteine P0DTC2
(b) The number of qubits that proteine P0DTC2 requires with one-hot and dense encoding
@@
(a) Exact and optimal value of proteine P0DTC2 on Qiskit
(b) Error rate of P0DTC2 on Qiskit
Figure 4: Comparison of the exact and optimal value on proteine P0DTC2
@@
When representing the protein P0DTC2 using quantum qubits, we conducted a comparison between one-hot and dense encoding. In Figure 3(a), a comparison of the required number of qubits with one hot and dense encoding for every fragment with a fixed length of 8 on P0DTC2 can be computed by Python3 according to equations 2 and 1. The one-hot encoding requires a minimum of 16 qubits and goes up to 42 qubits for maximum representation. In contrast, dense encoding requires a minimum of 8 qubits and extends up to 21 qubits for maximum representation. Most fragments in one-hot and dense encoding utilize 14 and 28 qubits respectively. In Figure 3(b), we extended this comparison to assess the maximum number of qubits required for encoding the P0DTC2 protein sequence using both one-hot and dense encoding across fragments with lengths ranging from 6 to 20, which also from Python3 according to equations 2 and 1. These results underscore the substantial reduction in the number of qubits needed when using dense encoding, which is significantly less than half of the requirement for one-hot encoding. This indicates that dense encoding is a resource-efficient approach for mRNA codon optimization in quantum computing.
@@
The main results of our method on Qiskit and the exact values are visualized in Figure 4(a). In this figure, the red line represents the exact value for each fragment, while each data point is the value generated by our approach in quantum computing. A noteworthy observation is that the majority of data points align closely with the line $y=x$. This alignment indicates that the sequences produced by our proposed algorithm are close to the best sequences, illustrating the accuracy of our approach. Equation 4 represents the error rate $E_{rate}$ between the optimal value from the VQE algorithm and the exact value, where $V_{vqe}$ represents the optimal value from the VQE algorithm and $V_{exact}$ represents the exact value.
$E_{rate}=\frac{V_{vqe}-V_{exact}}{V_{exact}}$ (4)
In Figure 4(b), we present the error rate among all of the fragments, showing the proximity of data points to zero. This signifies the stability of our approach in optimizing mRNA codon. This also proves that our approach has the ability to complete mRNA codon optimization on quantum computers.

### v2

@@ 1 Introduction
In bioinformatics, computational problems often present significant challenges due to the massive amount of data and the substantial computational resources. Searching for effective solutions to these problems is important for the development of bioinformatics. Quantum computing emerges as a promising candidate due to its distinctive properties, including superposition, entanglement, and interference between computational paths. These properties enable quantum computers to process vast amounts of data in parallel, potentially overcoming the limitations of classical computing for bioinformatics tasks. Therefore, exploring quantum algorithms for bioinformatics holds significant promise for accelerating the advancements in this field.
@@ 2 Related research
In the realms of DNA or RNA vaccine development, bioinformatics leverages the analysis of extensive pathogen sequences to identify target antigenic proteins for mRNA sequence design. This expedites the vaccine design process, facilitating the creation of safe and effective vaccines tailored to combat specific viruses. The COVID-19 pandemic, which spread rapidly from 2019 to 2022, had a profound global impact, affecting a vast number of people. This surge strained healthcare systems, leading to a scarcity of health resources and thereby preventing many individuals from receiving timely treatment. This had wide-ranging consequences on global health and everyday life. Additionally, the pandemic triggered a surge in unemployment, trade disruptions, and a downturn in the global economy [4]. This underscored the urgent need to conduct research into vaccine development, particularly focusing on highly infectious viruses, to safeguard human health.
@@
  • mRNA vaccines - mRNA vaccines have emerged as a revolutionary approach to immunization, utilizing a fragment of messenger RNA (mRNA) to trigger the production of viral proteins (antigens) within the human body. These antigens, produced by human cells in response to the injected mRNA, stimulate the immune system to generate protective antibodies. This cutting-edge technique not only offers safety advantages by not using weakened or dead viruses but also exhibits rapid development, high protein expression efficiency, and shortened production timelines [5][7][8].
  • Structure of the mRNA sequence - As shown in Figure 2, the structure of mRNA sequences is composed of the following components: 5’ cap, 5’ Untranslated Region(5’ UTR), Open Reading Frame(ORF), 3’ Untranslated Region (UTR), and Ploy-A tail [9]. While the 5’ untranslated region (5’UTR) and the 3’ untranslated region (3’UTR) are essential in regulating protein expression, they do not directly determine the specific amino acids incorporated into the protein. It’s worth noting that this paper does not discuss the optimization of mRNA codons through the 5’UTR and 3’UTR regions. The ORF is a part of the DNA, RNA, or mRNA sequence, which begins with a start codon (usually AUG, CUG, UUG in mRNA) and ends with a stop codon(UAA, UAG, and UGA in mRNA). The ORF in mRNA serves as the blueprint for protein synthesis, dictating the specific sequence of amino acids to be incorporated with the resulting antigen protein. In addition, the ORF is crucial for the development of mRNA vaccines, as it also plays a pivotal role in protein expression and stability [10][11].
@@
    Traditional methods for solving this problem on classical computers can be computationally intensive, especially as the length of the target sequence increases. The complexity arises from the vast search space of possible codon combinations and the need to balance factors like codon usage bias and secondary mRNA structure. Deep learning, particularly neural networks, has shown promise in addressing complex bioinformatics problems, including mRNA codon optimization [13][14][15]. Neural networks can automatically learn meaningful representations of biological data. In the context of mRNA codon optimization, this means that the model can capture the essential features of codon sequences that lead to optimal protein expression. Pre-trained models on related tasks or datasets can be fine-tuned for specific mRNA codon optimization problems. This can leverage knowledge gained from one domain to enhance performance in another. While deep learning holds promise, it’s essential to note that the success of these models depends on the availability of high-quality training data and careful design of the neural network architecture. Additionally, interoperability and understanding the biological implications of the model’s predictions are crucial in the context of bioinformatics applications.
@@
      Quantum annealers, like those developed by D-Wave Systems, hold significant promise for solving combinatorial optimization problems [20][21][22]. Quantum annealing is utilized for practical applications, including in the field of mRNA codon optimization. [3] The general process of applying quantum annealer to combinatorial optimization problems, such as mRNA codon optimization involves several steps. One is problem definition, specifying the optimization goal and relevant constraints. The second is to transform the problem into an Ising model representation. Then, the quantum annealing devices are used to find the lowest energy state. The final step is to decode the result to obtain the mRNA codon sequence, reflecting an optimization solution for protein expression.
@@
      VQE is a quantum algorithm designed to find the minimum eigenvalue for a given Hamiltonian, making it well-suited for solving certain types of combinatorial problems [23]. Firstly, a combinatorial problem is defined using a Hamiltonian. Then, a circuit ansatz with variational parameters is optimized to the approximate ground state of the Hamiltonian corresponding to the minimum expectation value. These quantum circuits are executed on a quantum computer to obtain the expectation value of the Hamiltonian. The process is iterated until convergence or a satisfactory solution quality is reached.
@@
      QAOA is a variational quantum algorithm, belonging to a broader class of VQE. It is tailored to find approximate optimization problems’ solutions [24]. The algorithm involves mapping a specific problem to the Ising model. In the aspect of quantum circuit construction, quantum state preparation encodes potential solutions, and a quantum circuit with alternating layers of unitary operations is built. Parameters are adjusted using classical optimization techniques to minimize the expectation value of the Hamiltonian. Iterations continue until convergence is reached.
@@
We have developed an innovative quantum method that significantly reduced quantum computing resources with a dense way of encoding compared to the existing quantum approach [3]. However, our approach adds complexities to the Hamiltonian due to the multiple local terms. Quadratic unconstrained binary optimization (QUBO) and QAOA approach mandates a Hamiltonian composed of only two local terms and it needs to add extra qubits for the reduction of additional terms. But this reduction also results in an increase in the required number of qubits. To address these complexities, we have employed the VQE algorithm, as it offers a unique advantage: it does not demand the addition of qubits. VQE enables us to seek the ground energy of the objective function efficiently, following which we employ classical computing techniques to derive the optimal mRNA sequence based on the ground state.
@@
Qiskit platform – Our implementation of codon optimization is conducted using the Qiskit platform which is the gate circuit quantum simulator as a central tool for codon optimization. Nevertheless, current quantum computing resources are constrained by the number of qubits. For example, optimizing the full-length protein sequences would demand thousands of qubits, surpassing the capabilities of existing quantum hardware. We’ve adopted a pragmatic approach to overcome this limitation by dividing the protein sequence into smaller, more manageable fragments with shorter lengths. This strategic division enables us to operate with fewer qubits for each fragment, ensuring a more feasible and efficient optimization process. By presenting this innovative approach to mRNA codon optimization, we aim to contribute to the evolving field of quantum computing in bioinformatics, providing new perspectives and solutions to complex problems in the realm of genetic coding and sequence optimization. The Qiskit platform allows to execution of quantum circuits in two ways. The first one connects to a native QPU backend like superconducting qubits, taking the results from the latest quantum processors into account. The second one is using a quantum circuit simulator where the qubits are assumed to be perfect qubits. That intrinsically implies no decoherence of the qubits and that the quantum gates have no error. Error models can be applied on the simulator for realistic simulation, however, often become a barrier to quantum logic and algorithm development for future quantum applications. For this project, we employ noiseless quantum circuit simulation for the proof-of-concept development and demonstration of our algorithm, based on the Perfect Intermediate Scale Quantum (PISQ) [25] approach.
@@ 3 Method
Quantum computers, in particular, hold promise due to their distinctive quantum mechanical properties such as superposition, entanglement, and interference. These properties empower quantum computers to efficiently handle extensive quantum data sets, finding the optimal (lowest or highest) value in a vast space solution. Notably, Fox et al. [3] introduced a codon optimization algorithm implemented on quantum annealers and gate-based circuit quantum computers. This groundbreaking work offers a practical approach to mRNA sequence design, unlocking new possibilities in the realm of bioinformatics and genetic research. The research by these authors represents a crucial milestone in harnessing the capabilities of quantum computing for this innovative application.
@@ 3.1 Synonymous codons encoding
Two examples for encoding amino acids A and L are given in Table 2 using either the one-hot or the dense way. This encoding allows straightforward identification of the selected codon based on the binary pattern. However, it introduces redundant quantum states (explained in the Appendix), which in turn increases the complexity of the quantum system.
@@ 3.3 Quantum algorithm execution
The SamplingVQE on the Qiskit platform [26] was used to implement the mRNA codon optimization for the protein fragment with a length of 8 while utilizing between 8 and 21 qubits. Given the challenges in accessing real quantum computers with a sufficient number of qubits, simulators in Qiskit allow scientists to explore optimization problems in quantum computing. In addition, Python3 served to generate Hamiltonian and as an accessible programming language on Qiskit. In the final step, the outcome of Qiskit was translated to mRNA sequence through Python3.
@@ 4 Results
The protein P0DTC2 was an example of optimizing mRNA codon sequences. The P0DTC2 protein from the spike protein of 2019-nCoV, sourced from UniProt and consisting of 1273 amino acids, would require 2234 and 4417 qubits for dense and one-hot encoding respectively according to equations 1 and 2. However, the available number of qubits on current quantum computers is limited, with no quantum computer possessing more than 1000 qubits. It is unrealistic to achieve mRNA codon optimization for the full-length protein P0DTC2 on a quantum computer because of the large number of qubits required. To overcome the limitation, the protein sequence is divided into shorter fragments, and each fragment is optimized separately by VQE. This approach requires fewer qubits for each fragment, making the optimization more manageable than optimizing the whole protein sequence at once. We use paper [3] as our reference paper.
(a) The number of qubits that protein P0DTC2 requires with one-hot and dense encoding
(b) The required number of gates that protein P0DTC2 requires with one hot and dense encoding
@@
In Figure 3(a), we extended this comparison to assess the maximum number of qubits required for encoding the P0DTC2 protein sequence using both one-hot and dense encoding across fragments with lengths ranging from 6 to 20 according to equations 2 and 1. For each fragment with a length of 6, the maximum number of qubits required for dense encoding is 16, whereas one-hot encoding requires 32. On average, dense encoding uses 11 qubits while one-hot needs 21. For longer fragments with a length of 19, the average qubit count for dense encoding is 33 compared to 66 with one-hot encoding. The maximum number of qubits for dense is 40 while one-hot requires 80 qubits. These results underscore the substantial reduction in the number of qubits needed when using dense encoding, which is significantly less than half of the requirement for one-hot encoding. This indicates that dense encoding is a resource-efficient approach for mRNA codon optimization in quantum computing compared to the current method [3].
Figure 4: The required number of qubits distribution between one hot and dense encoding on protein P0DTC2
Figure 3(b) shows that the average number of quantum gates required per fragment for encoding on the P0DTC2 protein varies significantly depending on the encoding method - one-hot and dense. For short fragments with a length of 6, one-hot encoding requires 537 gates, whereas dense encoding only requires 303 gates. For fragments with a length of 19, 4000 gates are required by one-hot encoding while dense encoding requires approximately 3000 gates. It indicates that the dense method uses fewer gates than one-hot encoding.
In the paper, we divided the protein into 159 fragments, each with a length of 8 amino acids (excluding the first amino acid M, which has only one codon and does not need optimization). Consequently, the number of required qubits for each fragment falls within the range of 8 to 21, as determined by equation 2. Each fragment is used to construct a Hamiltonian via equation 6, and then Sampling VQE on Qiskit is employed to identify the optimal codon sequence for each fragment. For each fragment with a length of 8, we conducted a comparison of the required number of qubits between one-hot and dense encoding for the protein P0DTC2. In Figure 4, a comparison of the required number of qubits with one hot and dense encoding for every fragment with a fixed length of 8 on P0DTC2 can be computed by Python3 according to equations 2 and 1. The one-hot encoding requires a minimum of 16 qubits and goes up to 42 qubits for maximum representation. In contrast, dense encoding requires a minimum of 8 qubits and extends up to 21 qubits for maximum representation. Most fragments in one-hot and dense encoding utilize 14 and 28 qubits respectively.
Figure 5: Exact and optimal value of protein P0DTC2 on Qiskit
@@
The main results of our method on Qiskit and the exact values are visualized in Figure 5. In this figure, the red line represents the exact value for each fragment, while each data point is the value generated by our approach in quantum computing. A noteworthy observation is that the majority of data points align closely with the line $y=x$. This alignment indicates that the sequences produced by our proposed algorithm are close to the best sequences, illustrating the accuracy of our approach.
In conclusion, this significant reduction in the required number of qubits and gates with dense encoding suggests potential advantages over one-hot encoding from Figure 3(a) and 3(b). In Figure 5, the quantum approach presented in this paper is designed to identify the optimal mRNA sequence by ensuring the Hamiltonian value matches the exact value that can be calculated by Python3. These results show that the quantum method has the capability to optimize mRNA codons more effectively than the current one [3].